# Antithetic Noise in Diffusion Models

### v1

@@ 5.2.1 Estimating pixel-wise statistics
Across all statistics, both AMC and QMC achieve much shorter CIs than MC, with relative efficiencies ranging from 3.1 to 136. This strongly suggests our proposed method can dramatically reduce the number of samples needed for reliable uncertainty estimates. AMC and QMC estimators often yield comparable results. The performance of QMC depends on how the total budget is allocated between the size of the QMC point set and the number of random replicates, a trade-off we explore in B.3.
@@
AMC 0.35 (32.66) 0.39 (27.12) 0.23 (22.05) 0.04 (6.96)
CelebA MC 1.82 1.76 0.60 0.60
AMC 0.26 (50.16) 0.31 (33.15) 0.19 (10.18) 0.20 (8.27)
Church MC 1.67 1.67 1.02 0.70
AMC 0.14 (136.37) 0.16 (106.79) 0.20 (27.16) 0.40 (3.14)
5.2.2 Estimating reconstruction uncertainty in Bayesian inverse problems
Table 4: Reconstruction uncertainty in L1 and PSNR
Metric Group Mean CI Length Eff.
L1 AMC 21.86 0.83 1.54
MC 21.98 1.03
PSNR AMC 19.03 0.26 1.71
MC 18.98 0.34
In Bayesian inverse problems, where the goal is to recover a clean image from a corrupted observation, we aim to reliably estimate the expected reconstruction quality and uncertainty of a posterior sampling algorithm. In this work, we focus on DPS chung2023diffusion . Reconstruction error is measured using PSNR and L1 distance to true image. We use 200 human face images from the CelebA-HQ dataset as ground truth, and consider the inpainting task (details in Appendix B.3). For each corrupted image, we generate 50 DPS reconstructions using PN and RR noise pairs and compare their estimates. They both give consistent results, with AMC being $54\%-71\%$ more efficient, though with smaller gains than earlier cases. Other tasks are reported in Appendix B.3.
6 Potential, limitations, and future directions
We find that antithetic initial noise yields negatively correlated samples. This finding enables us to improve sample diversity and construct more accurate estimators. However, our method is not a panacea: the LPIPS diversity gains in Table 2 are consistent but moderate. Although we greatly improve pixel-wise statistic estimates, our method yields very small gains (often 1%) on neural-network-based metrics like MUSIQ and CLIP Score versus standard Monte Carlo, since the Transformers’ strong nonlinearity absorbs most of the diffusion outputs’ negative correlation.
@@ 6 Potential, limitations, and future directions
Looking forward, future directions include systematically testing the symmetry conjecture; integrating antithetic noise with existing noise-optimization methods to enhance performance; and further leveraging the strengths of QMC to improve sampling and estimation.
Acknowledgement
Guanyang Wang acknowledges support from the National Science Foundation through grant DMS–2210849 and an Adobe Data Science Research Award. Bowen Song and Liyue Shen acknowledge support from University of Michigan MICDE (Michigan Institute for Computational Discovery and Engineering) Catalyst Grant, and University of Michigan MIDAS (Michigan Institute for Data Science) PODS Grant.
@@ Appendix A Proofs
On the other hand, since $-Z\sim\mathbb{N}(0,I)$ given $Z\sim\mathbb{N}(0,1)$, we have $\operatorname{Var}(f(Z))=\operatorname{Var}(f(-Z))$ and then $\operatorname{Var}(f(Z)+f(-Z))=0$, as claimed. ∎
A.2 Proofs in Section 3.3
We first formally state and prove the claim for the univariate case in Section 3.3.
Proposition 3 (Univariate case).
[…]
@@
FKG inequlity guarantees $\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(f(Z),g(Z))\geq 0$ provided that $f,g$ are non-decreasing. Therefore, $\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(F(Z),F(-Z))=-\operatorname{Cov}_{Z\sim\mathbb{N}(0,1)}(F(Z),-F(-Z))\leq 0$. Since
@@
we have $\operatorname{Corr}\bigl{(}F(z),\,F(-z)\bigr{)}\leq 0$. ∎
Now we prove the general case:
Proof of Proposition 1.
Let $G\coloneq S\circ\text{DM}$. For each coordinate $j\in[d]$ fix a sign
$s_{j}=\begin{cases}+1,&\text{if }G\text{ is non--decreasing in }x_{j},\\ -1,&\text{if }G\text{ is non--increasing in }x_{j}.\end{cases}$
Similarly, define
The multivariate FKG inequality [10] for product Gaussian measures states that, when $U$ and $V$ are coordinate-wise non–decreasing, $\operatorname{Cov}\bigl{(}U(Z),V(Z)\bigr{)}\geq 0$. Apply it to the pair $\bigl{(}\widetilde{G},-\widetilde{H}\bigr{)}$: both components are non–decreasing, so
[…]
A.3 Proofs in Section 4
A.3.1 Monte Carlo estimator:

### v2

@@ 5.1 Pixel-wise statistics
Results: Across all statistics, both AMC and QMC have much shorter CIs than MC, with relative efficiencies ranging from 3.1 to 136. This strongly suggests our estimators can dramatically reduce cost for uncertainty estimates. AMC and QMC estimators yield comparable results. The performance of QMC depends on how the total budget is allocated between the size of the QMC point set and the number of random replicates, a trade-off we explore in further details in Appendix  B.6.
@@
AMC ($k=2$) 0.35 (32.66) 0.39 (27.12) 0.23 (22.05) 0.04 (9.73)
AMC ($k=8$) 0.36 (30.35) 0.39 (27.34) 0.24 (20.37) 0.03 (13.94)
CelebA MC 1.77 1.76 0.60 0.60
AMC ($k=2$) 0.26 (47.41) 0.31 (33.15) 0.19 (10.18) 0.20 (7.82)
AMC ($k=8$) 0.15 (130.69) 0.16 (115.79) 0.18 (10.90) 0.19 (8.49)
Church MC 1.66 1.64 1.02 0.82
AMC ($k=2$) 0.14 (134.13) 0.16 (103.80) 0.20 (27.16) 0.26 (9.84)
AMC ($k=8$) 0.15 (117.90) 0.16 (107.64) 0.19 (27.81) 0.25 (10.87)
Stable Diff. MC 3.86 3.95 2.18 4.03
QMC 1.36 (8.07) 1.43 (7.60) 1.03 (4.46) 1.88 (4.54)
AMC ($k=2$) 0.90 (18.53) 1.00 (15.53) 0.89 (6.25) 1.62 (6.15)
DiT MC 7.52 7.74 3.32 3.15
QMC 3.44 (4.78) 3.49 (4.91) 1.82 (3.31) 1.69 (3.46)
AMC ($k=2$) 3.13 (5.79) 3.12 (6.16) 1.72 (3.74) 1.74 (3.30)
5.2 Evaluating different diffusion inverse problem solvers
Bayesian inverse problems are important in medical imaging, remote sensing, and astronomy. Recently, many methods have been proposed that leverage diffusion priors to solve inverse problems. For instance, Zheng et al. (2025) surveys 14 such methods, yet this is still far from exhaustive. As a result, comprehensive benchmarking is costly. Meanwhile, most existing works on new diffusion inverse solvers (DIS) largely ignore uncertainty quantification, though it is critical for fair evaluation. This raises a key challenge: how to efficiently evaluate posterior sampling algorithms with reliable estimates of reconstruction quality and uncertainty. Our approach aims to efficiently address this gap.
We show that our antithetic estimator can save substantial computational cost for quantifying uncertainty by requiring far fewer samples than standard estimators. We focus on two popular DISs, Diffusion Posterior Sampling (DPS)  (Chung et al., 2023a) and Decomposed Diffusion Sampler (DDS)  (Chung et al., 2023b), and evaluate them on a range of tasks described below.
5.2.1 Human Face Reconstruction
We evaluate DPS across three inverse problems: inpainting, Gaussian deblurring, and super-resolution. Reconstruction error is measured using both PSNR and $L_{1}$ distance relative to the ground truth image. For all tasks, we use 200 human face images from the CelebA-HQ dataset. For each corrupted image, we generate 50 DPS reconstructions using both PN and RR noise pairs to compare their estimators. Operator-specific configurations (e.g., kernel sizes, mask ratios) are provided in Appendix B.6.
Results: As shown in Table 3, across all three tasks, AMC achieves substantially shorter confidence intervals than standard MC. This implies AMC yields efficiency gains, ranging from $54\%$$84\%$ in inpainting, $41\%$$54\%$ in super-resolution, and $34\%$$56\%$ in deblurring.
The reconstruction metrics ($L_{1}$, PSNR) differ by less than $0.2\%$ between the two estimators, showing that they produce consistent results. Meanwhile, our method offers a clear advantage in efficiency.
Table 3: Comparison of AMC vs. MC across tasks in DPS with efficiency $(\text{CI}_{\text{MC}}/\text{CI})^{2}$ in parentheses
L1 PSNR
Task AMC CI MC CI AMC CI MC CI
Inpainting 0.83 (1.54) 1.03 0.23 (1.84) 0.31
Super-resolution 1.01 (1.41) 1.20 0.24 (1.54) 0.30
Gaussian Deblur 0.89 (1.34) 1.03 0.23 (1.56) 0.29
5.2.2 Medical Image Reconstruction
We use the DDS (Chung et al., 2023b) as the reconstruction backbone and apply our antithetic noise initialization to estimate reconstruction confidence intervals. Given a single measurement, we generate 100 reconstruction pairs using 50-step DDS sampling. Reconstruction quality is evaluated against ground truth using $L_{1}$ distance and PSNR. Dataset details can be found in Appendix B.6.4.
Results: Both MC and AMC estimators give consistent estimates, with estimated $L_{1}$ error of $0.0194$ and $0.0195$, and PSNR values of $31.48$ and $31.45$, respectively. Similarly, we observe that AMC produces much shorter confidence intervals. The CI lengths decrease with relative efficiency of $1.44$ for $L_{1}$ and $1.37$ for PSNR. This shows that antithetic initialization consistently reduces estimator variance with no extra cost without degrading reconstruction fidelity in large-scale inverse problems.
Conclusion: These two experiments show that a single antithetic estimator reduces evaluation costs for diffusion inverse solvers by $34\%$$84\%$, a saving that becomes especially valuable given the large number of solvers in the literature. This makes large-scale benchmarking far more practical.
5.3 Diversity Improvement
Many diffusion tools generate a few images from one prompt to let users select a preferred one. However, these images may look similar with randomly sampled initial noise (Marwood et al., 2023; Sadat et al., 2024). We show that antithetic noise produces more diverse images than independent noise, without reducing image quality or increasing computational cost. Diversity is measured using pairwise SSIM (lower = more diverse) and LPIPS (Zhang et al., 2018) (higher = more diverse).
We evaluate the diversity metrics for both unconditional and conditional diffusion on image pairs generated following the same setup in Section 5.1. As shown in Table 4, antithetic noise pairs consistently lead to higher diversity, as indicated by lower SSIM and higher LPIPS scores. Importantly, image quality remains stable for both noise types. We expect our method can be easily integrated into existing diversity optimization techniques to further improve their performance.
Metric Unconditional Diffusion Conditional Diffusion
CIFAR-10 LSUN-Church CelebA-HQ SD1.5 DiT
SSIM (%) 88.78 45.69 36.78 28.32 23.99
LPIPS (%) 6.69 3.54 15.14 5.78 10.62
6 Discussion and future directions
We find that antithetic initial noise yields negatively correlated samples. This is a robust property that generalizes across every model that we have tested. The negative correlation is especially useful for uncertainty quantification, where it brings huge variance reduction and cost savings. This makes our approach highly effective for evaluation and benchmarking tasks. Meanwhile, this finding can serve as a simple plug-in to improve related tasks, such as diversity improvement and image editing.
A limitation of our work is that while the symmetry conjecture is supported by both empirical and theoretical evidence, it remains open in full generality. In addition, our method is not a cure-all: the LPIPS diversity improvements in Table 4 are consistent but modest. While we obtain large gains for pixel-wise statistics, the improvements for perceptual or text-alignment metrics such as MUSIQ and CLIP Score are small, since the strong nonlinearity of transformer decoders attenuates most of the negative correlation present in the diffusion outputs.
@@ 6 Discussion and future directions
Looking forward, future directions include systematically studying the symmetry conjecture; integrating antithetic noise with existing noise-optimization methods; and further leveraging QMC to improve sampling and estimation.
@@ Appendix A Proofs
On the other hand, since $Z\sim\mathbb{N}(0,I)$ given $Z\sim\mathbb{N}(0,1)$, we have $\operatorname{Var}(f(Z))=\operatorname{Var}(f(-Z))$ and then $\operatorname{Var}(f(Z)+f(-Z))=0$, as claimed. ∎
A.2 Theoretical guarantee of the symmetry conjecture for large $t$
Background: The Ornstein–Uhlenbeck (OU) process.
The OU process $(X_{t})_{t\geq 0}$ in $\mathbb{R}^{d}$ solves the SDE
$\mathrm{d}X_{t}=-X_{t}\,\mathrm{d}t+\sqrt{2}\,\mathrm{d}B_{t},\qquad X_{0}\sim\mu_{0},$
where $(B_{t})$ is standard Brownian motion in $\mathbb{R}^{d}$. This is a Gaussian Markov process with stationary distribution $\gamma_{d}=\mathbb{N}(0,I_{d})$ and generator
$Lf(x)=\Delta f(x)-x\cdot\nabla f(x).$
For any initial law $\mu_{0}$, let $\mu_{t}$ denote the law of $X_{t}$. Then $\mu_{t}=\mu_{0}P_{t}$, where $(P_{t})_{t\geq 0}$ is the OU semigroup defined by
$(P_{t}f)(x)=\mathbb{E}[f(X_{t})\mid X_{0}=x].$
Write the score of $\mu_{t}$ as $s_{t}(x):=\nabla\log p_{t}(x)$, where $p_{t}$ is Lebesgue density of $\mu_{t}$, we have
Theorem 1 (Score converges to the Gaussian score).
Under the setup above,
$\lim_{t\to\infty}\ \mathbb{E}_{\mu_{t}}\!\bigl[\|s_{t}(X_{t})+X_{t}\|^{2}\bigr]\;=\;0.$
Moreover, the convergence is quantified by
$\mathbb{E}_{\mu_{t}}\!\bigl[\|s_{t}(X_{t})+X_{t}\|^{2}\bigr]\;=\;\mathcal{I}(\mu_{t}\mid\gamma_{d})\;\leq\;e^{-2t}\,\mathcal{I}(\mu_{0}\mid\gamma_{d}).$
@@ A.2 Theoretical guarantee of the symmetry conjecture for large $t$
Recall the definition of relative Fisher information
$\mathcal{I}(\mu_{t}\mid\gamma_{d}):=\mathbb{E}_{\mu_{t}}\!\bigl[\|\nabla\log(p_{t}/\gamma_{d})\|^{2}\bigr].$
Since $\nabla\gamma_{d}(x)=-x$, we have:
$\mathcal{I}(\mu_{t}\mid\gamma_{d})=\mathbb{E}_{\mu_{t}}\!\bigl[\|s_{t}(X_{t})+X_{t}\|^{2}\bigr].$
Moreover, it is well known (e.g. Chapter 5 of Bakry et al. (2013)) that $\mathcal{I}(\mu_{t}\mid\gamma_{d})\leq\exp(-2t)\mathcal{I}(\mu_{0}\mid\gamma_{d})$, this concludes our proof. ∎
In variance-preserving diffusion models, the forward noising process is exactly the Ornstein–Uhlenbeck semigroup. Theorem 1 implies that the true score converges to the Gaussian score x-x. Consequently, one-dimensional slices of the score become asymptotically affine antisymmetric, confirming our symmetry conjecture in the high-noise limit.
The next corollary shows, the 1-step DDIM update has nearly 1-1 correlation at large $t$.
Corollary 1 (Correlation of 1-step DDIM).
With all the setup the same as above, assuming we have a score network $\epsilon^{(t)}$ approximating the score function. Let $\eta_{t}:=\max\{\mathbb{E}_{p_{t}}\|\epsilon^{(t)}(X)-s_{t}(X)\|^{2},\mathbb{E}_{p_{t}}\|\epsilon^{(t)}(-X)-s_{t}(-X)\|^{2}\}$ be the expected (symmetrized) error. Consider one-step DDIM update $F_{t}(x)=a_{t}x+b_{t}\,\epsilon^{(t)}_{\theta}(x):=(F_{t,1}(x),\ldots,F_{t,d}(x))\in\mathbb{R}^{d}$ as defined in (2). Let
$v_{t,i}:=\min\{\operatorname{Var}_{p_{t}}\!\big(F_{t,i}(X)\big),\operatorname{Var}_{p_{t}}\!\big(F_{t,i}(-X)\big)\}.$
Then
@@
$\bigl|\operatorname{Corr}\bigl(F_{t,i}(X),F_{t,i}(-X)\bigr)+1\bigr|\;\leq\;\frac{2\,|b_{t}|}{\sqrt{v_{t,i}}}\Bigl(\sqrt{\eta_{t}}+e^{-t}\sqrt{\mathcal{I}(\mu_{0}\mid\gamma_{d})}\Bigr).$
Proof.
Theorem 1 shows $\mathcal{I}(\mu_{t}\mid\gamma_{d})\leq\exp(-2t)\mathcal{I}(\mu_{0}\mid\gamma_{d})$. Write $\epsilon^{(t)}_{\theta}=s_{t}+r_{t}$ and $s_{t}=-x+\Delta_{t}$, so that $\mathbb{E}_{p_{t}}\|r_{t}(X)\|^{2}\leq\eta_{t}$ and $\mathbb{E}_{p_{t}}\|\Delta_{t}(X)\|^{2}=\mathcal{I}(\mu_{t}\mid\gamma_{d})$. Then
$F_{t}(x)=(a_{t}-b_{t})x+b_{t}\bigl(\Delta_{t}(x)+r_{t}(x)\bigr),\qquad F_{t}(-x)=-F_{t}(x)+E_{t}(x),$
with $E_{t}(x):=b_{t}\!\left[\Delta_{t}(-x)+\Delta_{t}(x)+r_{t}(-x)+r_{t}(x)\right]$. By triangle inequality of $L^{2}(p_{t})$ norm,
$\|E_{t}\|_{L^{2}(p_{t})}\;\leq\;|b_{t}|\bigl(e^{-t}\sqrt{\mathcal{I}(\mu_{0}\mid\gamma_{d})}+e^{-t}\sqrt{\mathcal{I}(\mu_{0}\mid\gamma_{d})}+2\sqrt{\eta_{t}}\bigr).$
Using
$\operatorname{Corr}\bigl(F_{t,i}(X),F_{t,i}(-X)\bigr)=-1+\frac{\operatorname{Cov}\bigl(F_{t,i}(X),E_{t,i}(X)\bigr)}{\sqrt{\operatorname{Var}(F_{t,i}(X))\,\operatorname{Var}(F_{t,i}(-X))}}$
and Cauchy–Schwarz with the variance lower bounds gives
$\bigl|\operatorname{Corr}\bigl(F_{t}(X),F_{t}(-X)\bigr)+1\bigr|\;\leq\;\frac{\|E_{t}\|_{L^{2}(p_{t})}}{\sqrt{v_{t,i}}}\;\leq\;\frac{2\,|b_{t}|}{\sqrt{v_{t,i}}}\bigl(\sqrt{\eta_{t}}+e^{-t}\sqrt{\mathcal{I}(\mu_{0}\mid\gamma_{d})},$
as claimed. ∎
In practice, if $v_{t,i}>0$, the correlation is close to 1-1 for large $t$; its deviation from 1-1 is on the order of the neural network approximation error plus an exponentially decaying term $Ce^{-t}$.
A.3 Monotone convergence in $t$
Section A.2 shows that the score converges to the Gaussian score as $t\to\infty$. We now further prove that (1) the score error $\mathbb{E}_{\mu_{t}}\bigl[|s_{t}(X_{t})+X_{t}|^{2}\bigr]$ and (2) the density ratio $p_{t}/\gamma_{d}$ both converge monotonically in $t$.
A.3.1 Density ratio convergence
Background: Hermite Polynomials in $\mathbb{R}$:
For $x\in\mathbb{R}$, the (probabilist’s) Hermite polynomials $\{H_{n}(x)\}_{n\geq 0}$ form a sequence of polynomials defined by
$H_{n}(x)=(-1)^{n}\exp\left(\frac{x^{2}}{2}\right)\frac{\mathrm{d}^{n}}{\mathrm{d}x^{n}}\exp\left(-\frac{x^{2}}{2}\right).$
We summarize their known properties here. The first two can be checked by direct calculation. The latter two can be found on page 105 of Bakry et al. (2013).
Proposition 1.
Hermite polynomials $\{H_{n}(x)\}_{n\geq 0}$ satisfy the following
  • (Recurrence relation) $H_{n+1}(x)=x\,H_{n}(x)-H_{n}^{\prime}(x),\qquad n\geq 0.$
  • (Orthogonality under Gaussian measure) We have:
    $\int H_{n}(x)H_{m}(x)\gamma_{1}(\mathrm{d}x)=n!\delta_{nm},$
    where $\gamma_{1}(\mathrm{d}x)=(2\pi)^{-1/2}e^{-x^{2}/2}\mathrm{d}x$ denotes the one-dimensional standard Gaussian measure, and $\delta_{nm}$ is $1$ when $n=m$ and 0 otherwise.
  • (Completeness) Hermite polynomials $\{H_{n}(x)\}_{n\geq 0}$ form an orthogonal basis of the Hilbert space $L^{2}(\gamma):=\{f:\int f^{2}\gamma(\mathrm{d}x)<\infty\}$ equipped with inner product $\langle f,g\rangle_{\gamma}:=\int fg\gamma(\mathrm{d}x)$.
  • (Eigenfunction) Let $L$ be a differential operator defined as $L(f):=f^{\prime\prime}-xf^{\prime}$. Hermite polynomials $\{H_{n}(x)\}_{n\geq 0}$ are eigenfunctions of $L$:
    $L(H_{n})=-nH_{n},\qquad n\geq 0.$
Hermite Polynomials in $\mathbb{R}^{d}$: One can naturally extend univariate Hermite polynomials to the multivariate setting. For $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$, let $\gamma_{d}$ denote the $d$-dimensional standard Gaussian measure,
$\gamma_{d}(\mathrm{d}x):=(2\pi)^{-d/2}\exp\!\left(-\frac{\|x\|^{2}}{2}\right)\mathrm{d}x.$
For a multi-index $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}^{d}$, we define the multivariate Hermite polynomial
$H_{\alpha}(x):=\prod_{i=1}^{d}H_{\alpha_{i}}(x_{i}),$
where $H_{\alpha_{i}}$ is the one-dimensional probabilists’ Hermite polynomial. We write $|\alpha|:=\alpha_{1}+\cdots+\alpha_{d}$ and $\alpha!:=\alpha_{1}!\cdots\alpha_{d}!$. Similar to the one-dimensional case, the family $\{H_{\alpha}\}_{\alpha\in\mathcal{N}^{d}}$ forms an orthogonal basis of the Hilbert space $L^{2}(\gamma_{d})$:
$\int_{\mathbb{R}^{d}}H_{\alpha}(x)\,H_{\beta}(x)\,\gamma_{d}(\mathrm{d}x)=\alpha!\,\delta_{\alpha\beta}.$
Monotone convergence of the density ratio.
Recall from Section A.2 that the generator of the OU process
$Lf(x)=\Delta f(x)-x\cdot\nabla f(x).$
and the corresponding Markov semigroup $P_{t}=\exp(tL)$. We show that the multivariate Hermite polynomials are eigenfunctions of $L$, and thus also of $P_{t}$.
Proposition 2.
We have:
  • $LH_{\alpha}=-\,|\alpha|\,H_{\alpha},\qquad\alpha\in\mathbb{N}^{d}$,
  • $P_{t}H_{\alpha}=e^{-\,|\alpha|t\,}H_{\alpha},\qquad\alpha\in\mathbb{N}^{d},t\geq 0$.
  • For every $g\in L^{2}(\gamma_{d})$
    $P_{t}g(x)=\sum_{\alpha\in\mathbb{N}^{d}}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x),$
    where $a_{\alpha}=\langle g,H_{\alpha}\rangle_{L^{2}(\gamma_{d})}$.
Proof.
We can directly calculate
$LH_{\alpha}(x)$ $=\Delta H_{\alpha}(x)-x\cdot\nabla H_{\alpha}(x)$
$=\sum_{i}H_{\alpha_{i}}^{\prime\prime}(x_{i})H_{\alpha_{-i}}(x_{-i})-\sum_{i}x_{i}H_{\alpha_{i}}^{\prime}(x_{i})H_{\alpha_{-i}}(x_{-i})$
$=\sum_{i}H_{\alpha_{-i}}(x_{-i})(H_{\alpha_{i}}^{\prime\prime}(x_{i})-x_{i}H_{\alpha_{i}}^{\prime}(x_{i}))$
$=\sum_{i}(-\alpha_{i})H_{\alpha_{-i}}(x_{-i})H_{\alpha_{i}}(x_{i})$
$=-\,|\alpha|\,H_{\alpha}(x),$
where $H_{\alpha_{-i}}(x_{-i})=\prod_{j\neq i}H_{\alpha_{j}}(x_{j})$. The second to last equality follows from the fourth property of Proposition 1.
Consequently, the OU semigroup $(P_{t})_{t\geq 0}$ satisfies
$P_{t}H_{\alpha}=e^{-t|\alpha|}H_{\alpha},\qquad t\geq 0.$
For every $g\in L^{2}(\gamma_{d})$, we rewrite $g$ according to its Hermite expansion (which is valid according to properties 2-3 in Proposition 1).
$g(x)=\sum_{\alpha\in\mathbb{N}^{d}}a_{\alpha}H_{\alpha}(x),\qquad a_{\alpha}=\langle g,H_{\alpha}\rangle_{L^{2}(\gamma_{d})}.$
Therefore
$P_{t}g(x)=\sum_{\alpha\in\mathbb{N}^{d}}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x),$
as claimed. ∎
In particular, let $p_{t}(x)=f_{t}(x)\gamma_{d}(x)$ denote the density at time $t$ of the OU process started from $p_{0}$. It is known that $f_{t}=P_{t}f_{0}$. Proposition 2 implies
$f_{t}(x)=\sum_{\alpha\in\mathbb{N}^{d}}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x)=a_{0}+\sum_{|\alpha|>0}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x),$
where $a_{\alpha}$ are the Hermite coefficients of $f_{0}$. Since $H_{0}\equiv 1$ and
$a_{0}=\int_{\mathbb{R}^{d}}f_{0}(x)\,H_{0}(x)\,\gamma_{d}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}f_{0}(x)\,\gamma_{d}(\mathrm{d}x)=\int_{\mathbb{R}^{d}}p_{0}(x)\,\mathrm{d}x=1,$
we may rewrite
$f_{t}(x)=1+\sum_{|\alpha|>0}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x).$ (3)
We now show $f_{t}$ monotonically converges to $1$ in $L^{2}(\gamma_{d})$.
Proposition 3.
Let $p_{t}=f_{t}\gamma_{d}$ be as above, and assume $f_{0}\in L^{2}(\gamma_{d})$. Then $\|f_{t}-1\|_{L^{2}(\gamma_{d})}$ monotonically decays to 0 at the speed of $e^{-t}$.
Proof.
From (3) we have
$f_{t}(x)=1+\sum_{|\alpha|>0}a_{\alpha}e^{-t|\alpha|}H_{\alpha}(x).$
Since $\{H_{\alpha}\}_{\alpha\in\mathbb{N}^{d}}$ is an orthogonal basis of $L^{2}(\gamma_{d})$, we obtain
$\|f_{t}-1\|_{L^{2}(\gamma_{d})}^{2}=\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{d}\end{subarray}}a_{\alpha}^{2}e^{-2t|\alpha|}.$
Since each term in the sum decreases monotonically to 0, the same holds for $\|f_{t}-1\|_{L^{2}(\gamma_{d})}^{2}$. The overall decay rate is governed by the slowest mode, which corresponds to the smallest $\lvert a\rvert=1$, thus the rate of convergence is $\exp(-2t)$. Taking square roots yields the desired result. ∎
A.3.2 Score error convergence
Recall from Section A.2 that $\mathbb{E}_{\mu_{t}}\!\bigl[\|s_{t}(X_{t})+X_{t}\|^{2}\bigr]=\mathcal{I}(\mu_{t}\mid\gamma_{d})$. It follows from equation 5.7.2 in Bakry et al. (2013) that
$\frac{\mathrm{d}\mathcal{I}(\mu_{t}\mid\gamma_{d})}{\mathrm{d}t}\leq-2\mathcal{I}(\mu_{t}\mid\gamma_{d})\leq 0.$
Thus $\mathcal{I}(\mu_{t}\mid\gamma_{d})$ is monotonically decreasing as a function of $t$. Further, Grönwall’s inequality implies
$\mathcal{I}(\mu_{t}\mid\gamma_{d})\;\leq\;e^{-2t}\,\mathcal{I}(\mu_{0}\mid\gamma_{d}).$
A.4 Symmetry preservation under the forward process
This section shows the forward process preserves any orthogonal symmetry in the data distribution. More precisely, let $G$ be a subgroup of the orthogonal group $O(d)$ in $\mathbb{R}^{d}$ (e.g., coordinate sign flips, coordinate permutations, and rotations). We have the following:
Proposition 4.
Assuming $p_{0}$ is symmetric about $\mu$ under the action of $G$, i.e., for every $g\in G$, we have $p_{0}(x)=p_{0}(g\cdot(x-\mu)+\mu)$ for every $x$. Then $p_{t}$ satisfies $p_{t}(x)=p_{t}(g\cdot(x-\mu_{t})+\mu_{t})$ for every $x$, and $s_{t}(\mu_{t}+g\cdot x)=g\cdot s_{t}(\mu_{t}+x)$, where $\mu_{t}=e^{-t}\mu$.
Proof.
Let $X_{0}\sim p_{0},Z\sim\mathbb{N}(0,I_{d})$. It is known that $X_{t}:=e^{-t}X_{0}+\sqrt{1-e^{-2t}}Z\sim p_{t}$. For any $g\in G$,
$g\cdot(X_{t}-\mu_{t})+\mu_{t}$ $=g\cdot(e^{-t}X_{0}+\sqrt{1-e^{-2t}}Z-\mu_{t})+\mu_{t}$
$=e^{-t}(g\cdot(X_{0}-\mu)+\mu)+\sqrt{1-e^{-2t}}g\cdot Z\sim p_{t}$
where the last claim uses that $g\cdot(X_{0}-\mu)+\mu$ has the same distribution as $X_{0}$ (by symmetry) and $g\cdot Z\sim\mathbb{N}(0,I_{d})$. This proves the claim on the symmetry of $p_{t}$. For the score, since $p_{t}(\mu_{t}+g\cdot x)=p_{t}(\mu_{t}+x)$, taking derivative on both sides yields:
$g^{\top}\cdot\nabla p_{t}(\mu_{t}+g\cdot x)=\nabla p_{t}(\mu_{t}+x).$
Since $g$ is an orthogonal matrix, $g^{\top}=g^{-1}$, we have:
$\nabla p_{t}(\mu_{t}+g\cdot x)=g\cdot\nabla p_{t}(\mu_{t}+x).$
Dividing both sides by $p_{t}(\mu_{t}+x)$ proves the claimed result. ∎
Sections A.2 (large $t$ limit), A.3 (monotone convergence), and A.4 (symmetry perservation from $t=0$) together provide insights for the conjecture: The forward OU process preserves all orthogonal symmetries of $p_{0}$, including central reflections. Hence, in the idealized case where $p_{0}$ is reflection-symmetric, the density is even and its score is odd. Proposition 4 ensures that this symmetry is preserved for all $t$. Furthermore, even when the symmetry is only approximate near $t\approx 0$, the monotone convergence of $p_{t}$ toward the Gaussian is exponentially fast, so the forward dynamics push $s_{t}$ toward a linear function in a monotone manner.
A.5 Proofs in Section 4
A.5.1 Monte Carlo estimator:
Proposition 5.